There are three consecutive integers listed in natural order. the sum of the second and thrid is half the first decreased by 6.
HELP! Thanks

Answer by ReadingBoosters(3246)

x x+1 x+2
x+1 + x+2 = 1/2x-6
2x + 3 = 1/2x - 6
2x - 1/2x = -6 - 3
1.5x = -9
x = -6
Integers
-6, -5, -4
Proof
-5 + -4 = -9
1/2(-6)=-3 - 6 = -9